OpenStudy (anonymous):

4│x + 5│ - 2 ≤ 10 How do you solve this ???

OpenStudy (anonymous):

well , these lines ----> | xxx | mean absolute value so you should know what to do there :]

OpenStudy (helder_edwin):

use this theorem: if \(b\geq0\), \[ \large |a|\leq b\quad\Leftrightarrow -b\leq a\leq b \]

jhonyy9 (jhonyy9):

so like the first step i think you can assuming to the both sides 2 and like step 2. so i think you will can divide both sides by 4 and after that check what will get hope so much that you can calculi these two steps so wait your reply with this work good luck